OpenStudy (anonymous):

HELP ME!.!.!.!..... What is the common difference of the following arithmetic sequence? 102, 100, 98, 96…

OpenStudy (anonymous):

:) yay help meh

OpenStudy (anonymous):

96-98=? 98-100=? 100-102=?

OpenStudy (anonymous):

O_e !?!?!

OpenStudy (anonymous):

What is 96-98?

OpenStudy (anonymous):

dont no

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

You dont know what 96-98 is?

OpenStudy (anonymous):

\[96-98\] \[= -2\]

OpenStudy (anonymous):

OMG i thought it was 102

OpenStudy (anonymous):

Umm, no it's -2

OpenStudy (anonymous):

What is 98-100?